### v1

# Convergence, Duality and Well-Posedness in Convex Bilevel Optimization

Abstract
We consider the convex bilevel optimization problem, also known as simple bilevel programming. There are two challenges in solving convex bilevel optimization problems. Firstly, strong duality is not guaranteed due to the lack of Slater constraint qualification. Secondly, we demonstrate through an example that convergence of algorithms is not guaranteed even when usual subotimality gap bounds are present, due to the possibility of encountering super-optimal solutions. We show that strong duality (but not necessarily dual solvability) is exactly equivalent to ensuring correct asymptotic convergence of both inner and outer function values, and provide a simple condition that guarantees strong duality. Unfortunately, we also show that this simple condition is not sufficient to guarantee convergence to the optimal solution set. We draw connections to Levitin-Polyak well-posedness, and leverage this together with our strong duality equivalence to provide another condition that ensures convergence to the optimal solution set. We also discuss how our conditions have been implicitly present in existing algorithmic work.
keywords:
convex bilevel optimization; simple bilevel programming; duality; well-posedness
{MSCcodes}
90C25; 90C30; 90C46
1 Introduction
Consider a constrained convex optimization problem given by
(1) $g^{*}=\min_{x\in X}g(x),$
where $X$ is a finite-dimensional closed convex set and $g:X\to{\mathbb{R}}$ is a convex function. Here, $g^{*}$ is the optimal value of Eq. 1. In general, problem (1) might have multiple optimal solutions, and in this case it is natural to consider a second objective to select the best solution among the optimal solutions with respect to an additional criterion. Supposing that this is given by a convex function $f:{X}\to\mathbb{R}$, this leads to the following so-called convex bilevel optimization problem (also known as simple bilevel program):
(2) $f^{*}=\min_{x\in X_{g}}f(x),\quad\text{where }X_{g}=\operatorname*{arg\,min}_{z\in X}g(z).$
Here, $f^{*}$ denotes the optimal value of the convex bilevel problem Eq. 2, and $X_{g}$ is the set of minimizers for Eq. 1. Throughout this paper, we assume that problems (1) and (2) are solvable. Moreover, we denote an optimal solution of (2) by $x^{*}$, satisfying $f(x^{*})=f^{*}$ and $g(x^{*})=g^{*}$.
In contrast to a standard single-level convex optimization problem, the presence of two objective functions in the convex bilevel problem (2) brings in additional challenges, which we categorize into two groups. First, we do not have an explicit representation of the optimal set $X_{g}$ in general. This rules out the possibility of applying projection-based algorithms as well as the conditional gradient method, making both the projection onto $X_{g}$ and the linear optimization oracle over $X_{g}$ intractable. With this first challenge in mind, a stream of literature in this domain has alternatively considered the value function formulation of (2):
(3) $\min_{x\in X}\left\{f(x):\ g(x)\leq g^{*}\right\}.$
This is now a convex optimization problem with a convex functional constraint. However, a critical issue, i.e., our second challenge, is that (3) does not satisfy strict feasibility and hence the standard Slater constraint qualification condition fails. As a result, it is possible that the Lagrangian dual of (3) is unsolvable, and in such a case (3) will violate a critical assumption exploited in the convergence analysis of most primal-dual methods. One may attempt to enforce Slater’s condition by adding a small $\epsilon_{g}>0$ to the right-hand side of the constraint, i.e., $g(x)\leq g^{*}+\epsilon_{g}$, but this approach provides guarantees only in terms of $\epsilon_{g}$-optimality and does not solve the actual problem (3). In fact, it was demonstrated in [16, Appendix D] that such an approach is numerically unstable.
1.1 Contributions
Approaches to solve single-level optimization problems cannot by default be applied to Eq. 2 (or equivalently Eq. 3), and thus it requires specific algorithms for the bilevel setting, with different analyses. In this paper, we examine the algorithmic guarantees which exist for various bilevel optimization algorithms, and provide conditions to reassure us of convergence to an optimum solution of Eq. 2. Algorithmic guarantees for Eq. 2 generally ensure that an iteratively computed sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfies the following suboptimality gap bounds
(4) $f(x_{t})\leq f^{*}+r_{f,t},\quad g(x_{t})\leq g^{*}+r_{g,t},\quad\textbf{where}\lim_{t\to\infty}r_{f,t}=\lim_{t\to\infty}r_{g,t}=0.$
Unfortunately, as we will demonstrate in Example 2.1, this is not sufficient to ensure that $\{x_{t}\}_{t\geq 1}$ converges to a solution of Eq. 2. Our study aims to analyze this carefully, and provide conditions that ensure convergence. We examine two notions of convergence: (i) convergence in objective function value Eq. 5; and (ii) convergence in distance to the optimal set Eq. 7.
Our main result establishes a previously unknown equivalence: strong duality (but not necessarily dual solvability) for the value-function formulation (3) is equivalent to ensuring convergence in objective function value. Despite the lack of Slater condition, we present two simple conditions that ensure strong duality: one based on a proximity assumption for $g$ (Section 3), and the other based on compactness of the optimal set (Section 3). While the proximity assumption is natural, it unfortunately does not ensure convergence in distance to the optimal set, as Example 3.3 demonstrates. On the other hand, we show that the compactness assumption is equivalent to two notions of Levitin-Polyak well-posedness of Eq. 2, which yields both convergence in objective value as well as in distance to the optimal set. Our strong duality equivalence is further used to provide quantifiable bounds on the distance. These findings provide algorithm-independent criteria under which iterate sequences for Eq. 2 have meaningful convergence guarantees.
1.2 Related work
Much of the work on convex bilevel optimization involves the design and analysis of algorithms to solve Eq. 2. There are several different schemes, including iterative regularization [3, 26, 27, 14, 18, 9, 1, 17, 24, 13, 20, 22, 6, 25], sublevel set methods [2, 16, 4, 10, 5, 28] and sequential averaging [21, 23, 19].
Our work focuses the duality theory of Eq. 2, namely the Lagrange dual of its value function formulation Eq. 3. Friedlander and Tseng [12] show that correctness of the exact regularization technique is equivalent to the existence of an optimal dual multiplier of Eq. 3, and they demonstrate a number of special cases where exact regularization holds, e.g., conic programs and exact penalty methods for convex functional constraints. Dempe et al. [8], Franke et al. [11] provide conditions to guarantee existence of an optimal dual multiplier by using more general constraint qualifications than Slater. In contrast, our work focuses on when strong duality holds regardless of whether the dual is solvable or not, and in particular we highlight that this case has important algorithmic implications for Eq. 2. Consequently, we are able to provide weaker conditions for strong duality, and thus convergence assurance, than the ones above that guarantee that an optimal dual multiplier exists. Friedlander and Tseng [12] observe that, under coercivity assumptions, strong duality holds (without solvability), which is a simple consequence of Sion’s minimax theorem, though they did not explore this further.
Finally, we note that there is a broader literature on general bilevel optimization problems with two sets of variables, which are non-convex; see Dempe [7] for details. This literature also examines optimality conditions via Lagrange multipliers, though the theory is different to the simpler convex problem Eq. 2 that we consider.
1.3 Notation and basic assumption
We let $\mathbb{R}$} and $\mathbb{N}$} denote the real and natural numbers, respectively. Given any positive integer $T$, let $T]:=\{1,2,\dots,T\}$. For any $a\in{\mathbb{R}}$, we let $a]_{+}:=\max\{0,a\}$. We define $\{a_{t}\}_{t\in[T]}$ to be the sequence of elements $(a_{1},a_{2},\dots,a_{T})$. Let $\|\cdot\|_{2}$ be the Euclidean norm. For a convex function $f:X\to{\mathbb{R}}$, we denote the gradient at $x\in X$ by $\nabla f(x)$.
We will assume that Eq. 2 satisfies the following properties throughout the paper. {assumption} The set $X$ is closed and convex, and belongs to some finite-dimensional Euclidean space. The functions $f$ and $g$ are convex and real-valued on all of $X$. Furthermore, the minimization problem Eq. 1 is solvable, i.e., $g^{*}<\infty$ and $X_{g}$ defined in Eq. 2 is non-empty and closed. Finally, the bilevel optimization problem Eq. 2 is solvable, and we denote the set of optimal solutions to Eq. 2 as
$X^{*}:=\operatorname*{arg\,min}_{x\in X_{g}}f(x)=\{x\in X_{g}:f(x)=f^{*}\},$
which is non-empty and closed.
2 Super-optimality and Lagrange duality
The central issue in solving Eq. 2 is that, in practice, many algorithms very rarely find a feasible solution $\bar{x}\in X_{g}$. Instead, algorithms will generate a point $x_{t}\in X$ with $g(x_{t})>g^{*}$. Since $X_{g}\subset\left\{x\in X:g(x)\leq g(x_{t})\right\}$, the point $x_{t}$ may be super-optimal for Eq. 2, meaning that $f(x_{t})<f^{*}$. As iterative algorithms progress, the accuracy of $x_{t}$ (hopefully) improves, i.e., it is shown that $g(x_{t})\leq g^{*}+r_{g,t}$ where $r_{g,t}\geq 0$ are error terms such that $r_{g,t}\to 0$. Since $g(x_{t})\geq g^{*}$, this implies that $\lim_{t\to\infty}g(x_{t})=g^{*}$. For the outer-level objective $f$, it is often similarly shown that $f(x_{t})\leq f^{*}+r_{f,t}$, where $r_{f,t}\geq 0$ are error terms such that $r_{f,t}\to 0$. Since we might have $f(x_{t})<f^{*}$, these inequalities are only sufficient to deduce
(5) $\limsup_{t\to\infty}f(x_{t})\leq f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*},$
but they do not guarantee $f(x_{t})\to f^{*}$. Indeed, we may have a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ for which $\underline{f}:=\lim_{k\to\infty}f(x_{t_{k}})<f^{*}$. On the other hand, it is straightforward to show that any limit point of $\{x_{t}\}_{t\geq 1}$ is an optimal solution of Eq. 2. That said, it is difficult to infer in an online fashion whether the current $f(x_{t})$ is close to $f^{*}$ or $\underline{f}$. In other words, we have no assurance of the approximation quality of $f(x_{t})$.
We now present an example that demonstrates how examining $f(x_{t})$ without further assurances can indeed be misleading for estimating $f^{*}$.
Example 2.1.
Suppose that
[…]
$g(x)$ $=\frac{x_{2}^{2}}{x_{1}}$
$X$ $=\left\{x\in{\mathbb{R}}^{2}:\,x_{1}\geq 1,~0\leq x_{2}\leq 1\right\}.$
Clearly, $f,g,X$ are convex. Furthermore, we have $X_{g}=\{(s,0)\in{\mathbb{R}}^{2}:\,s\geq 1\}$, $g^{*}=0$, $X^{*}=\{(s,0)\in{\mathbb{R}}^{2}:s\geq 100\}$ and $f^{*}=0$. However, consider now the sequence $\{x_{t}\}_{t\geq 1}\subset X$ given by $x_{t}:=(t,1)$ for $t\geq 1$. Then $g(x_{t})=1/t\to 0=g^{*}$, but $f(x_{t})=[100-t]_{+}^{2}-1\to-1<f^{*}$. Thus, $\{x_{t}\}_{t\geq 1}$ satisfies Eq. 5 but $f(x_{t})$ is not a good estimate for $f^{*}$.
To guard against situations like in Example 2.1, and recognizing the numerous algorithmic works which provide guarantees of the form Eq. 5, we look for conditions which ensure that whenever Eq. 5 holds, the following guarantee also holds:
(7) $\lim_{t\to\infty}f(x_{t})=f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*}.$
Another guarantee that is of interest is that the distance to the optimal solution set of Eq. 2 decreases to 0:
(8) $\lim_{t\to\infty}\operatorname{Dist}(x_{t},X^{*})=0,\quad\text{where }\operatorname{Dist}(x,X^{*}):=\inf_{z\in X^{*}}\|x-z\|_{2}.$
Interestingly, we uncover a relationship between Lagrange duality and Eq. 7. The Lagrange dual of Eq. 2 is obtained via Eq. 3, and is given by
(9) $d^{*}:=\sup_{\lambda\geq 0}q(\lambda),\quad\text{where}\quad q(\lambda):=\inf_{x\in X}\left\{f(x)+\lambda(g(x)-g^{*})\right\}.$
Here, $\lambda\geq 0$ is the dual variable, $q(\cdot)$ is the (concave) dual objective function, and $d^{*}$ is the optimal dual value. Since $g(x)\geq g^{*}$ for any $x\in X$, we can deduce that $q(\lambda)$ is non-decreasing in $\lambda$. This means that $d^{*}=\sup_{\lambda\geq 0}q(\lambda)=\lim_{\lambda\to\infty}q(\lambda)$. Recalling that $f^{*}$ is the optimal value of Eq. 3, by weak duality we always have $d^{*}\leq f^{*}$. We will use the following definition of strong duality.
Definition $2.$
We say that strong duality holds for Eq. 2 when $d^{*}=f^{*}$.
Note that Definition 2.2 does not guarantee that the dual is solvable, but only that $q(\lambda)\to f^{*}$ as $\lambda\to\infty$. Therefore, ensuring that strong duality holds requires much weaker conditions than ensuring that the dual is solvable, as was studied in previous works [12, 8, 11].
Our main result is that ensuring convergence Eq. 7 is exactly equivalent to strong duality holding for Eq. 2.
Theorem 2.3.
Consider Eq. 2 when $f,g,X$ are convex. Then the following are equivalent:
  1. 1.
    Any sequence $\{x_{t}\}_{t\geq 1}\subset X$ that satisfies Eq. 5 also satisfies Eq. 7.
  2. 2.
    Strong duality (Definition 2.2) holds for Eq. 2.
We will use the following lemma in the proof of Theorem 2.3, as well as subsequent results.
Lemma 2.4.
There exists a sequence $\{x_{t}\}_{t\geq 1}\subset X$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$.
Proof 2.5 (Proof of Lemma 2.4).
We consider two cases: $d^{*}>$-$\infty$ and $d^{*}=$-$\infty$. In the first case, there must exist some $\lambda_{0}\geq 0$ such that $q(\lambda_{0})=\inf_{x\in X}\left\{f(x)+\lambda_{0}(g(x)-g^{*})\right\}>$-$\infty$, and consequently, $q(\lambda)\geq q(\lambda_{0})>$-$\infty$ for any $\lambda\geq\lambda_{0}$. For each $t\geq 1$, choose $x_{t}\in X$ to be a point that satisfies $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})+1/t$. By definition of $q(\cdot)$, we also have $f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\geq q(\lambda_{0})$, so combining with $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})$ we get
$g(x_{t})-g^{*}\leq\frac{1}{t}\left(q(t+\lambda_{0})+1/t-q(\lambda_{0})\right).$
This implies that $g(x_{t})\to g^{*}$ as $t\to\infty$. Now, from the choice of $x_{t}\in X$, using the definition of $g_{*}$ we have $g(x_{t})\geq g_{*}$ and also we have
$f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})$ $\leq f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})$
$\leq q(t+\lambda_{0})+1/t.$
Taking limit superior of all terms in the inequality, we have
$\limsup_{t\to\infty}f(x_{t})=\limsup_{t\to\infty}\left(f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\right)\leq\limsup_{t\to\infty}\left(q(t+\lambda_{0})+1/t\right)=d^{*}.$
The first equality follows because $g(x_{t})\to g^{*}$, and the last equality follows since $\lim_{t\to\infty}q(t+\lambda_{0})=d^{*}$ (recall the definition of $d_{*}$ and the fact that $q(\lambda)$ is non-decreasing in $\lambda$).
Now we consider the second case when $d^{*}=$-$\infty$. Then for $t\geq 1$, let $z_{t}\in X$ be a point such that $f(z_{t})+t(g(z_{t})-g^{*})\leq-t$, and define $\theta_{t}:=\frac{1}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\in[0,1]$. Let $x^{*}\in X^{*}$ be a solution to Eq. 2, and define $x_{t}:=\theta_{t}z_{t}+(1-\theta_{t})x^{*}\in X$. Notice that $g(x_{t})\leq\theta_{t}g(z_{t})+(1-\theta_{t})g(x^{*})=g^{*}+\theta_{t}(g(z_{t})-g^{*})=g^{*}+\frac{g(z_{t})-g^{*}}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\leq g^{*}+1/\sqrt{t}\to g^{*}$. Thus, $\lim_{t\to\infty}g(x_{t})=g^{*}$. On the other hand, $f(x_{t})\leq\theta_{t}f(z_{t})+(1-\theta_{t})f(x^{*})$, and by definition of $z_{t}$, we have $f(z_{t})\leq-t(1+g(z_{t})-g^{*})$, so $\theta_{t}f(z_{t})\leq-\frac{\sqrt{t}(1+g(z_{t})-g^{*})}{\max\{1,g(z_{t})-g^{*}\}}$. But this implies $\theta_{t}f(z_{t})\to$-$\infty$ as $t\to\infty$, therefore $\limsup_{t\to\infty}f(x_{t})=-\infty=d^{*}$.
Proof 2.6 (Proof of Theorem 2.3).
Suppose that strong duality holds, i.e., $f_{*}=d_{*}$. Let $\{x_{t}\}_{t\geq 1}$ be a sequence satisfying Eq. 5. Then given any $\epsilon>0$, there exists $\lambda_{\epsilon}\in[0,\infty)$ such that
$\inf_{x\in X}\left\{f(x)+\lambda_{\epsilon}(g(x)-g^{*})\right\}\geq f^{*}-\epsilon.$
This implies that for any $t\geq 1$ we have $f(x_{t})\geq f^{*}-\lambda_{\epsilon}(g(z_{t})-g^{*})-\epsilon$. By taking limit inferior as $t\to\infty$ on both sides, and recognizing that $\lim_{t\to\infty}g(x_{t})=g^{*}$, we obtain $\liminf_{t\to\infty}f(x_{t})\geq f^{*}-\epsilon$. Since $\epsilon>0$ is chosen arbitrarily, $\liminf_{t\to\infty}f(z_{t})\geq f^{*}$, which implies (7).
Suppose now that strong duality does not hold, so $d^{*}=\sup_{\lambda\geq 0}q(\lambda)<f^{*}$. Lemma 2.4 states that there exists a sequence $\{x_{t}\}_{t\geq 1}$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Since $d^{*}<f^{*}$, Eq. 5 holds for $\{x_{t}\}_{t\geq 1}$, but not Eq. 7.
3 Conditions for strong duality and convergence
We now provide two conditions to ensure that strong duality holds for Eq. 2, thereby (equivalently) ensuring that any algorithm which guarantees Eq. 5 has assurance that Eq. 7 holds also. {condition} If $\{x_{t}\}_{t\geq 1}\subset X$ is a sequence such that $g(x_{t})\to g^{*}$, we have $\operatorname{Dist}(x_{t},X_{g})\to 0$.
{condition}
The set of optimal solutions $X^{*}$ for Eq. 2 is compact.
Section 3 generalizes, for example, error bound-type conditions on $g$, where $\tau\operatorname{Dist}(x,X_{g})^{r}\leq g(x)-g^{*}$. Error bounds are used in several works [e.g., 16, 5, 20] to enable stronger guarantees for algorithms solving Eq. 2.
Section 3 can be shown to be equivalent to the coercivity of the functions $h=\max\{f,g\}$ and $\bar{h}=\max\{f-f^{*},g-g^{*}\}$ over $X$. (Recall that a function $h$ is coercive over $X$ if $h(x)\to\infty$ along any recessive direction in $X$.) Section 3, or more restrictive versions of it, has appeared in various works. Cabot [3], Solodov [26, 27], Malitsky [18], Dempe et al. [9] assume exactly Section 3. Helou and Simões [14], Amini and Yousefian [1], Kaushik and Yousefian [17], Shen et al. [24], Giang-Tran et al. [13], Jiang et al. [16], Cao et al. [5] employ the assumption that $X$ is bounded to prove guarantees for their algorithms. This immediately implies Section 3 holds. Finally, Beck and Sabach [2], Sabach and Shtern [21], Shehu et al. [23], Doron and Shtern [10], Merchav and Sabach [19] use the assumption that $f$ itself is coercive, which also implies Section 3. The coercivity assumption on $f$ is commonly motivated by learning problems where $g$ models the prediction error and $f$ is a regularization function, e.g., the $\ell_{1}$- or $\ell_{2}$-norm.
Section 3 guarantees that strong duality holds, and hence Eq. 5 implies Eq. 7.
Proposition 3.1.
Suppose that Section 3 holds for Eq. 2. Then strong duality holds, and, consequently, any sequence satisfying Eq. 5 also satisfies Eq. 7.
Proof 3.2 (Proof of Proposition 3.1).
Suppose that $\{x_{t}\}_{t\geq 1}$ is a sequence which satisfies the premise of Lemma 2.4: $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Let $x^{*}\in X^{*}$ be an optimal solution to Eq. 2, and define the sequence $\{y_{t}\}_{t\geq 1}\subset X_{g}$ by $y_{t}:=\operatorname*{arg\,min}_{y\in X_{g}}\|x_{t}-y\|_{2}$. (Note that the minimizer is unique since the $\ell_{2}$-norm is strictly convex.) By convexity of $f$, we have
$f(x_{t})-f^{*}$ $=f(x_{t})-f(x^{*})$
$\geq\nabla f(x^{*})^{\top}(x_{t}-x^{*})$
$=\nabla f(x^{*})^{\top}(x_{t}-y_{t})+\nabla f(x^{*})^{\top}(y_{t}-x^{*})$
$\geq-\|\nabla f(x^{*})\|_{2}\|x_{t}-y_{t}\|_{2}+\nabla f(x^{*})^{\top}(y_{t}-x^{*})$
$\geq-\|\nabla f(x^{*})\|_{2}\|x_{t}-y_{t}\|_{2}.$
The second inequality follows from the Cauchy-Schwarz inequality, while the third follows since $\nabla f(x^{*})^{\top}(y_{t}-x^{*})\geq 0$ by the optimality of $x^{*}$ in minimizing $f$ over $X_{g}$, and the fact that $y_{t}\in X_{g}$. By Section 3, we have $\|x_{t}-y_{t}\|_{2}=\operatorname{Dist}(x_{t},X_{g})\to 0$ since $g(x_{t})\to g^{*}$. Therefore, the limit inferior of both sides is
$\liminf_{t\to\infty}f(x_{t})-f^{*}\geq 0.$
In summary, we have
$f^{*}\leq\liminf_{t\to\infty}f(x_{t})\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*},$
so strong duality holds.
Unfortunately, Section 3 on its own is not enough to guarantee that Eq. 8 holds.
Example 3.3.
Define $X=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1\}\subset{\mathbb{R}}^{3}$, and $g(x)=x_{3}^{2}$. Then $g^{*}=0$ and $X_{g}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1,x_{3}=0\}$. Notice that for any $x\in X$, $\operatorname{Dist}(x,X_{g})=|x_{3}|=\sqrt{g(x)}$, therefore Section 3 is satisfied. Now define $f(x)=x_{2}^{2}/x_{1}$, which is convex. Then $f^{*}=0$ and $X^{*}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,x_{2}=0,x_{3}=0\}$. Now consider the sequence $\{x_{t}\}_{t\geq 1}$ defined by $x_{t}=(t,1,0)\in X_{g}$. Then $f(x_{t})=1/t\to f^{*}$, $g(x_{t})=0=g^{*}$, $\operatorname{Dist}(x_{t},X_{g})=0$ but $\operatorname{Dist}(x_{t},X^{*})=1$ for all $t$. Therefore, Eq. 8 is not satisfied.
Section 3 is more powerful than Section 3, and in fact has strong connections to the study of well-posedness, which examines properties of approximation schemes for solutions of optimization (and related) problems.
Definition 3.4 (Huang and Yang [15]).
We say that Eq. 2 satisfies:
  • Levitin-Polyak (LP) well-posedness if, for any sequence $\{x_{t}\}_{t\geq 1}\subset X$ which satisfies $\operatorname{Dist}(x_{t},X_{g})\to 0$ and $f(x_{t})\to f^{*}$, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
  • generalized LP well-posedness if, for any sequence that satisfies Eq. 7, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
  • strong generalized LP well-posedness if, for any sequence that satisfies Eq. 5, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
It is immediately clear that strong generalized LP well-posedness implies generalized LP well-posedness. We will show that every sense of well-posedness in Definition 3.4 implies Section 3. It turns out that strong generalized LP well-posedness is equivalent to Section 3 in finite dimensions (which is our setting), which follows from a result of Huang and Yang [15]. We will additionally show that Section 3 ensures that strong duality holds, and thus Theorem 2.3 implies that generalized LP is equivalent to strong generalized LP well-posedness for Eq. 2.
Proposition 3.5.
The following are equivalent for Eq. 2:
  1. (a)
    Section 3 holds.
  2. (b)
    Eq. 2 is generalized LP well-posed.
  3. (c)
    Eq. 2 is strong generalized LP well-posed.
Furthermore, if Eq. 2 is LP well-posed, then Section 3 holds.
Proof 3.6 (Proof of Proposition 3.5).
First, note that if $X^{*}$ is not compact, then it must be unbounded, since we assume that $X^{*}$ is closed. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X^{*}$ such that $\|x_{t}\|_{2}\to\infty$. Then this sequence satisfies all three premises of Definition 3.4, however there is no convergent subsequence. Therefore all three senses of well-posedness imply compactness of $X^{*}$, i.e., Section 3. Huang and Yang [15, Theorem 2.4] states that when $X$ is finite-dimensional, which we assume, then strong generalized LP well-posedness is equivalent to Section 3.
To show that strong generalized LP is equivalent to generalized LP, we will first show that under Section 3, strong duality holds. First, denote
$X_{1}:=\left\{x\in X:~f(x)-f^{*}\leq 1,~g(x)-g^{*}\leq 1\right\}.$
Note that $X_{1}$ is the $1$-sublevel set of the function $\bar{h}=\max\{f-f^{*},g-g^{*}\}$, and $X^{*}$ is the 0-sublevel set. Since $X^{*}$ is bounded, all sublevel sets of $\bar{h}$ are bounded, hence $X_{1}$ is also bounded. Let $\{x_{t}\}_{t\geq 1}\subset X$ be a sequence satisfying the premise of Lemma 2.4, i.e., $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. By weak duality $d^{*}\leq f^{*}$, therefore $f(x_{t})\leq f^{*}+1$ eventually, and also $g(x_{t})\leq g^{*}+1$ eventually as well. This means that only finitely many $x_{t}\not\in X_{1}$, and since $X_{1}$ is bounded, the whole sequence is bounded. Take a convergent subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$. Since $X$ is closed, $\bar{x}\in X$. Furthermore, since $g$ is convex, it is lower semicontinuous, so $g^{*}=\liminf_{k\to\infty}g(x_{t_{k}})\geq g(\bar{x})\geq g^{*}$, hence $\bar{x}\in X_{g}$. But this implies that $f(\bar{x})\geq f^{*}$ by definition of $f^{*}$. Therefore $f^{*}\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*}$ (where the last inequality follows from the premise on the sequence $\{x_{t}\}_{t\geq 1}$), and thus strong duality holds.
Now, assume that Eq. 2 is generalized LP well-posed. Then $X^{*}$ must be compact. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfying Eq. 5. Since compactness of $X^{*}$ implies strong duality, from Theorem 2.3 we know that $\{x_{t}\}_{t\geq 1}$ also satisfies Eq. 7. Thus, by generalized LP well-posedness, there exists $\bar{x}\in X^{*}$ and a subsequence for which $x_{t_{k}}\to\bar{x}$, hence Eq. 2 is also strong generalized LP well-posed. Since a strong generalized LP well-posed problem is also generalized LP well-posed, this shows the equivalence.
This equivalence yields a quantitative version of Eq. 8.
Corollary 3.7.
Suppose that Section 3 holds for Eq. 2. Then there exists a function $c:{\mathbb{R}}_{+}^{2}\to{\mathbb{R}}_{+}$ such that $c(0,0)=0$ and for any $x\in X$,
$|f(x)-f^{*}|\geq c\left(\operatorname{Dist}(x,X^{*}),g(x)-g^{*}\right).$
Furthermore, the function $c$ satisfies that whenever $\beta_{t}\to 0$, $c(\alpha_{t},\beta_{t})\to 0$, then $\alpha_{t}\to 0$. Therefore, under Section 3 any sequence $\{x_{t}\}_{t\geq 1}$ satisfying Eq. 5 also satisfies Eq. 8.
Proof 3.8 (Proof of Corollary 3.7).
Proposition 3.5 states that Section 3 is equivalent to generalized LP well-posedness. Huang and Yang [15, Theorem 2.2] then ensures that whenever Eq. 2 is generalized LP well-posed, the function $c$ that satisfies the required properties is
$c(\alpha,\beta):=\inf_{x}\left\{|f(x)-f^{*}|:~x\in X,~\operatorname{Dist}(x,X^{*})=\alpha,~g(x)-g^{*}=\beta\right\}.$