### v2

# Convergence, Duality and Well-Posedness in Convex Bilevel Optimization

Abstract
We consider the convex bilevel optimization problem, also known as simple bilevel programming. Strong duality of the convex bilevel optimization problem is not guaranteed, due to the lack of Slater constraint qualification. Furthermore, convergence of algorithms is in general not guaranteed without further conditions due to possibility of super-optimal solutions. We show that strong duality (but not necessarily dual solvability) is exactly equivalent to ensuring correct asymptotic convergence of function values. We provide a simple condition that guarantees strong duality. Unfortunately, we show that this is not sufficient to guarantee convergence to the optimal solution set. We draw connections to Levitin-Polyak well-posedness, and leverage this together with our strong duality equivalence to provide another condition that ensures convergence to the optimal solution set with quantifiable bounds.
1 Introduction
Consider a constrained convex optimization problem given by
$g^{*}=\min_{x\in X}g(x),$ (1)
where $X$ is a finite-dimensional closed convex set and $g:X\to{\mathbb{R}}$ is a convex function. Here, $g^{*}$ is the optimal value of Eq. 1. In general, problem (1) might have multiple optimal solutions, and in this case it is natural to consider a second objective to select the best solution among the optimal solutions with respect to an additional criterion.
Supposing that this is given by a convex function $f:{X}\to\mathbb{R}$, this leads to the following so-called convex bilevel optimization problem (also known as simple bilevel program):
$f^{*}=\min_{x\in X_{g}}f(x),\quad\text{where }X_{g}=\operatorname*{arg\,min}_{z\in X}g(z).$ (2)
Here, $f^{*}$ denotes the optimal value of the convex bilevel problem Eq. 2, and $X_{g}$ is the set of minimizers for Eq. 1. Throughout this paper, we assume that problems (1) and (2) are solvable. Moreover, we denote an optimal solution of (2) by $x^{*}$, satisfying $f(x^{*})=f^{*}$ and $g(x^{*})=g^{*}$.
In contrast to a standard single-level convex optimization problem, the presence of two objective functions in the convex bilevel problem (2) brings in additional challenges, which we categorize into two groups. First, we do not have an explicit representation of the optimal set $X_{g}$ in general. This rules out the possibility of applying projection-based algorithms as well as the conditional gradient method, making both the projection onto $X_{g}$ and the linear optimization oracle over $X_{g}$ intractable.
With this first challenge in mind, a stream of literature in this domain has alternatively considered the value function formulation of (2):
$\min_{x\in X}\left\{f(x):\ g(x)\leq g^{*}\right\}.$ (3)
This is now a convex optimization problem with a convex functional constraint. However, a critical issue, i.e., our second challenge, is that (3) does not satisfy strict feasibility and hence the standard Slater constraint qualification condition fails. As a result, it is possible that the Lagrangian dual of (3) is unsolvable, and in such a case (3) will violate a critical assumption exploited in the convergence analysis of most primal-dual methods. One may attempt to enforce Slater’s condition by adding a small $\epsilon_{g}>0$ to the right-hand side of the constraint, i.e., $g(x)\leq g^{*}+\epsilon_{g}$, but this approach provides guarantees only in terms of $\epsilon_{g}$-optimality and does not solve the actual problem (3). In fact, it was demonstrated in [16, Appendix D] that such an approach is numerically unstable.
Contributions.
Approaches to solve single-level optimization problems cannot by default be applied to Eq. 2 (or equivalently Eq. 3), and thus it requires specific algorithms for the bilevel setting, with different analyses. In this short communication, we examine the algorithmic guarantees which exist for various bilevel optimization algorithms, and provide conditions to reassure us of convergence. Algorithmic guarantees for Eq. 2 generally ensure that an iteratively computed sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfies the following inequalities:
$f(x_{t})\leq f^{*}+r_{f,t},\quad g(x_{t})\leq g^{*}+r_{g,t},\quad\textbf{where}\lim_{t\to\infty}r_{f,t}=\lim_{t\to\infty}r_{g,t}=0.$ (4)
Unfortunately, as we will demonstrate in Section 3, this is not sufficient to ensure that $\{x_{t}\}_{t\geq 1}$ converges to a solution of Eq. 2. Our study aims to analyze this carefully, and provide conditions which ensure convergence. It turns out that this can be studied by examining the regularity of the Lagrange dual of Eq. 3. As mentioned above, Eq. 3 does not satisfy Slater’s condition is by definition. That said, we can still analyze its Lagrange dual, and examine conditions for strong duality to hold. Here, we must differentiate between two cases: (i) one where the optimal dual value is equal to $f^{*}$; and a stronger case where additionally there is an optimal dual multiplier $\lambda^{*}$. We will show that the first case is in fact equivalent to ensuring that sequences satisfying Eq. 4 will in fact converge to an optimal solution of Eq. 2. This uncovers a useful hidden regularity structure in Eq. 2 that, to the best of our knowledge, has not been elaborated on previously. We conclude the paper by presenting two conditions to ensure that strong duality holds, and examine how these conditions have been implicitly present in existing algorithmic work.
Related work.
Much of the work on convex bilevel optimization involves the design and analysis of algorithms to solve Eq. 2. There are several different schemes, including iterative regularization [3, 26, 27, 14, 18, 9, 1, 17, 24, 13, 20, 22, 6, 25], sublevel set methods [2, 16, 4, 10, 5, 28] and sequential averaging [21, 23, 19].
Our work focuses on the duality theory of Eq. 2, namely the Lagrange dual of its value function formulation Eq. 3. Friedlander and Tseng [12] show that correctness of the exact regularization technique is equivalent to the existence of an optimal dual multiplier of Eq. 3, and they demonstrate a number of special cases where exact regularization holds, e.g., conic programs and exact penalty methods for convex functional constraints. Dempe et al. [8], Franke et al. [11] provide conditions to guarantee existence of an optimal dual multiplier by using more general constraint qualifications than Slater. In contrast, our work focuses on when strong duality holds regardless of whether the dual is solvable or not, and in particular we highlight that this case has important algorithmic implications for Eq. 2. Consequently, we are able to provide weaker conditions for strong duality, and thus convergence assurance, than the ones above that guarantee that an optimal dual multiplier exists. Friedlander and Tseng [12] observe that, under boundedness assumptions, strong duality holds (without solvability), which is a simple consequence of Sion’s minimax theorem, though they did not explore this further.
Finally, we note that there is a broader literature on general bilevel optimization problems with two sets of variables, which are non-convex; see Dempe [7] for details. This literature also examines optimality conditions via Lagrange multipliers, though the theory is different to the simpler problem Eq. 2 that we consider.
Notation and basic assumption.
We let $\mathbb{R}$} and $\mathbb{N}$} denote the real and natural numbers, respectively. Given any positive integer $T$, let $T]:=\{1,2,\dots,T\}$. For any $a\in{\mathbb{R}}$, we let $a]_{+}:=\max\{0,a\}$. We define $\{a_{t}\}_{t\in[T]}$ to be the sequence of elements $(a_{1},a_{2},\dots,a_{T})$. Let $\|\cdot\|_{2}$ be the Euclidean norm. For a convex function $f:X\to{\mathbb{R}}$, we denote the gradient at $x\in X$ by $\nabla f(x)$. We will assume that Eq. 2 satisfies the following properties throughout the paper.
Assumption 1.
The set $X$ is closed and convex, and belongs to some finite-dimensional Euclidean space. The functions $f$ and $g$ are convex and real-valued on all of $X$. Furthermore, the minimization problem Eq. 1 is solvable, i.e., $g^{*}<\infty$ and $X_{g}$ defined in Eq. 2 is non-empty and closed. Finally, the bilevel optimization problem Eq. 2 is solvable, and we denote the set of optimal solutions to Eq. 2 as
$X^{*}:=\operatorname*{arg\,min}_{x\in X_{g}}f(x)=\{x\in X_{g}:f(x)=f^{*}\},$
which is non-empty and closed.
2 Super-optimality and Lagrange duality
The central issue in solving Eq. 2 is, in practice, we very rarely encounter a feasible solution $\bar{x}\in X_{g}$. Instead, algorithms will generate a point $x_{t}\in X$ with $g(x_{t})>g^{*}$. Since $X_{g}\subset\left\{x\in X:g(x)\leq g(x_{t})\right\}$, the point $x_{t}$ may be super-optimal for Eq. 2, meaning that $f(x_{t})<f^{*}$. As iterative algorithms progress, the accuracy of $x_{t}$ (hopefully) improves, i.e., it is shown that $g(x_{t})\leq g^{*}+r_{g,t}$ where $r_{g,t}\geq 0$ are error terms such that $r_{g,t}\to 0$. Since $g(x_{t})\geq g^{*}$, this implies that $\lim_{t\to\infty}g(x_{t})=g^{*}$. For the outer-level objective $f$, it is often similarly shown that $f(x_{t})\leq f^{*}+r_{f,t}$, where $r_{f,t}\geq 0$ are error terms such that $r_{f,t}\to 0$. Since we might have $f(x_{t})<f^{*}$, these inequalities are only sufficient to deduce
$\limsup_{t\to\infty}f(x_{t})\leq f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*},$ (5)
but they do not guarantee $f(x_{t})\to f^{*}$. Indeed, we may have a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ for which $\underline{f}:=\lim_{k\to\infty}f(x_{t_{k}})<f^{*}$. On the other hand, it is straightforward to show that any limit point of $\{x_{t}\}_{t\geq 1}$ is an optimal solution of Eq. 2. That said, it is difficult to infer in an online fashion whether the current $f(x_{t})$ is close to $f^{*}$ or $\underline{f}$. In other words, we have no assurance of the approximation quality of $f(x_{t})$.
We now present an example that demonstrates how examining $f(x_{t})$ without further assurances can indeed be misleading for estimating $f^{*}$.
Example 1.
Suppose that
[…]
Clearly, $f,g,X$ are convex. Furthermore, we have $X_{g}=\{(s,0)\in{\mathbb{R}}^{2}:\,s\geq 1\}$, $g^{*}=0$, $X^{*}=\{(s,0)\in{\mathbb{R}}^{2}:s\geq 100\}$ and $f^{*}=0$. However, consider now the sequence $\{x_{t}\}_{t\geq 1}\subset X$ given by $x_{t}:=(t,1)$ for $t\geq 1$. Then $g(x_{t})=1/t\to 0=g^{*}$, but $f(x_{t})=[100-t]_{+}^{2}-1\to-1<f^{*}$. Thus, $\{x_{t}\}_{t\geq 1}$ satisfies Eq. 5 but $f(x_{t})$ is not a good estimate for $f^{*}$.
To guard against situations like in Example 1, and recognizing the numerous algorithmic works which provide guarantees of the form Eq. 5, we look for conditions which ensure that whenever Eq. 5 holds, the following also holds:
$\lim_{t\to\infty}f(x_{t})=f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*}.$ (7)
Another possible condition we consider is the that the distance to the optimal solution set of Eq. 2 decreases to 0:
$\lim_{t\to\infty}\operatorname{Dist}(x_{t},X^{*})=0,\quad\text{where }\operatorname{Dist}(x,X^{*}):=\inf_{z\in X^{*}}\|x-z\|_{2}.$ (8)
Interestingly, we uncover a relationship between Lagrange duality and Eq. 7. The Lagrange dual of Eq. 2 is obtained via Eq. 3, and is given by
$d^{*}:=\sup_{\lambda\geq 0}q(\lambda),\quad\text{where}\quad q(\lambda):=\inf_{x\in X}\left\{f(x)+\lambda(g(x)-g^{*})\right\}.$ (9)
Here, $\lambda\geq 0$ is the dual variable, $q(\cdot)$ is the (concave) dual objective function, and $d^{*}$ is the optimal dual value. Since $g(x)\geq g^{*}$ for any $x\in X$, we can deduce that $q(\lambda)$ is non-decreasing in $\lambda$. This means that $d^{*}=\sup_{\lambda\geq 0}q(\lambda)=\lim_{\lambda\to\infty}q(\lambda)$. Recalling that $f^{*}$ is the optimal value of Eq. 3, by weak duality we always have $d^{*}\leq f^{*}$. We will use the following definition of strong duality.
Definition 1.
We say that strong duality holds for Eq. 2 when $d^{*}=f^{*}$.
Note that Definition 1 does not guarantee that the dual is solvable, but only that $q(\lambda)\to f^{*}$ as $\lambda\to\infty$. Therefore, ensuring that strong duality holds requires much weaker conditions than ensuring that the dual is solvable, as was studied in previous works [12, 8, 11].
Our main result is that ensuring convergence Eq. 7 is exactly equivalent to strong duality holding for Eq. 2.
Theorem 1.
Considering Eq. 2 when $f,g,X$ are convex, the following are equivalent:
  1. 1.
    Any sequence $\{x_{t}\}_{t\geq 1}\subset X$ that satisfies Eq. 5 also satiesfies Eq. 7.
  2. 2.
    Strong duality (Definition 1) holds for Eq. 2.
We will use the following lemma in the proof of Theorem 1, as well as subsequent results.
Lemma 1.
There exists a sequence $\{x_{t}\}_{t\geq 1}\subset X$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$.
Proof of Lemma 1.
We consider two cases: $d^{*}>$-$\infty$ and $d^{*}=\infty$. In the first case, there must exist some $\lambda_{0}\geq 0$ such that $q(\lambda_{0})=\inf_{x\in X}\left\{f(x)+\lambda_{0}(g(x)-g^{*})\right\}>$-$\infty$, and consequently, $q(\lambda)\geq q(\lambda_{0})>$-$\infty$ for any $\lambda\geq\lambda_{0}$. For each $t\geq 1$, choose $x_{t}\in X$ to be a point that satisfies $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})+1/t$. We also have $f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\geq q(\lambda_{0})$, so combining with $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})$ we get
$g(x_{t})-g^{*}\leq\frac{1}{t}\left(q(t+\lambda_{0})+1/t-q(\lambda_{0})\right).$
This implies that $g(x_{t})\to g^{*}$. Now observe that
$f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})$ $\leq f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})$
$\leq q(t+\lambda_{0})+1/t.$
Taking limit superior of all terms in the inequality, we have
$\limsup_{t\to\infty}f(x_{t})=\limsup_{t\to\infty}\left(f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\right)\leq\limsup_{t\to\infty}\left(q(t+\lambda_{0})+1/t\right)=d^{*}.$
The first equality follows because $g(x_{t})\to g^{*}$, and the last equality follows since $\lim_{t\to\infty}=d^{*}$.
Now we consider the second case when $d^{*}=$-$\infty$. Then for $t\geq 1$, let $z_{t}\in X$ be a point such that $f(z_{t})+t(g(z_{t})-g^{*})\leq-t$, and define $\theta_{t}=\frac{1}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\in[0,1]$. Let $x^{*}\in X^{*}$ be a solution to Eq. 2, and define $x_{t}=\theta_{t}z_{t}+(1-\theta_{t})x^{*}\in X$. Notice that $g(x_{t})\leq\theta_{t}g(z_{t})+(1-\theta_{t})g(x^{*})=g^{*}+\theta_{t}(g(z_{t})-g^{*})=g^{*}+\frac{g(z_{t})-g^{*}}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\leq g^{*}+1/\sqrt{t}\to g^{*}$. Therefore $\lim_{t\to\infty}g(x_{t})=g^{*}$. On the other hand, $f(x_{t})\leq\theta_{t}f(z_{t})+(1-\theta_{t})f(x^{*})$, and by definition of $z_{t}$, we have $f(z_{t})\leq-t(1+g(z_{t})-g^{*})$, so $\theta_{t}f(z_{t})\leq-\frac{\sqrt{t}(1+g(z_{t})-g^{*})}{\max\{1,g(z_{t})-g^{*}\}}$. But this implies $\theta_{t}f(z_{t})\to$-$\infty$ as $t\to\infty$, therefore $\limsup_{t\to\infty}f(x_{t})=-\infty=d^{*}$. ∎
Proof of Theorem 1.
Suppose that strong duality holds. Let $\{x_{t}\}_{t\geq 1}$ be a sequence satisfying Eq. 5. Then given any $\epsilon>0$, there exists $\lambda_{\epsilon}\in[0,\infty)$ such that
$\inf_{x\in X}\left\{f(x)+\lambda_{\epsilon}(g(x)-g^{*})\right\}\geq f^{*}-\epsilon.$
This implies that for any $t\geq 1$ we have $f(x_{t})\geq f^{*}-\lambda_{\epsilon}(g(z_{t})-g^{*})-\epsilon$. By taking limit inferior as $t\to\infty$ on both sides, and recognizing that $\lim_{t\to\infty}g(x_{t})=g^{*}$, we obtain $\liminf_{t\to\infty}f(x_{t})\geq f^{*}-\epsilon$. Since $\epsilon>0$ is chosen arbitrarily, $\liminf_{t\to\infty}f(z_{t})\geq f^{*}$, which implies (7).
Suppose now that strong duality does not hold, so $d^{*}=\sup_{\lambda\geq 0}q(\lambda)<f^{*}$. Lemma 1 states that there exists a sequence $\{x_{t}\}_{t\geq 1}$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Since $d^{*}<f^{*}$, Eq. 5 holds for $\{x_{t}\}_{t\geq 1}$, but not Eq. 7. ∎
3 Conditions for strong duality and convergence
We now provide two conditions to ensure that strong duality holds for Eq. 2, thereby (equivalently) ensuring that any algorithm which guarantees Eq. 5 has assurance that Eq. 7 holds also.
Condition 1.
If $\{x_{t}\}_{t\geq 1}\subset X$ is a sequence such that $g(x_{t})\to g^{*}$, we have $\operatorname{Dist}(x_{t},X_{g})\to 0$.
Condition 2.
The set of optimal solutions $X^{*}$ for Eq. 2 is compact.
1 generalizes, for example, error bound-type conditions on $g$, where $\tau\operatorname{Dist}(x,X_{g})^{r}\leq g(x)-g^{*}$. Error bounds are used in several works [e.g., 16, 5, 20] to enable stronger guarantees for algorithms solving Eq. 2.
2 can be shown to be equivalent to coercivity of the functions $h=\max\{f,g\}$ and $\bar{h}=\max\{f-f^{*},g-g^{*}\}$. 2, or more restrictive versions of it, has appeared in various works. Cabot [3], Solodov [26, 27], Malitsky [18], Dempe et al. [9] assume exactly 2. Helou and Simões [14], Amini and Yousefian [1], Kaushik and Yousefian [17], Shen et al. [24], Giang-Tran et al. [13], Jiang et al. [16], Cao et al. [5] employ the assumption that $X$ is bounded to prove guarantees for their algorithms. This immediately implies 2 holds. Finally, Beck and Sabach [2], Sabach and Shtern [21], Shehu et al. [23], Doron and Shtern [10], Merchav and Sabach [19] use the assumption that $f$ is coercive, which also implies 2. This is commonly motivated by learning problems where $g$ models the prediction error and $f$ is a regularization function, e.g., the $\ell_{1}$- or $\ell_{2}$-norm.
1 guarantees that strong duality holds, and hence Eq. 5 implies Eq. 7.
Proposition 1.
Suppose that 1 holds for Eq. 2. Then strong duality holds, and, consequently, any sequence satisfying Eq. 5 also satisfies Eq. 7.
Proof of Proposition 1.
Suppose that $\{x_{t}\}_{t\geq 1}$ is a sequence satisfying the premise of Lemma 1: $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Let $x^{*}\in X^{*}$ be an optimal solution to Eq. 2, and define the sequence $\{y_{t}\}_{t\geq 1}\subset X_{g}$ by $y_{t}:=\operatorname*{arg\,min}_{y\in X_{g}}\|x_{t}-y\|_{2}$. (Note that the minimizer is unique since the $\ell_{2}$-norm is strictly convex.) By convexity of $f$, we have
$f(x_{t})-f^{*}$ $=f(x_{t})-f(x^{*})$
$\geq\nabla f(x^{*})^{\top}(x_{t}-x^{*})$
$=\nabla f(x^{*})^{\top}(x_{t}-y_{t})+\nabla f(x^{*})^{\top}(y_{t}-x^{*})$
$\geq-\|\nabla f(x^{*})\|_{2}\|x_{t}-y_{t}\|_{2}+\nabla f(x^{*})^{\top}(y_{t}-x^{*})$
$\geq-\|\nabla f(x^{*})\|_{2}\|x_{t}-y_{t}\|_{2}.$
The second inequality follows from the Cauchy-Schwarz inequality, while the third follows since $\nabla f(x^{*})^{\top}(y_{t}-x^{*})\geq 0$ by the optimality of $x^{*}$ in minimizing $f$ over $X_{g}$, and the fact that $y_{t}\in X_{g}$. By 1, we have $\|x_{t}-y_{t}\|_{2}=\operatorname{Dist}(x_{t},X_{g})\to 0$ since $g(x_{t})\to g^{*}$. Therefore, the limit inferior of both sides is
$\liminf_{t\to\infty}f(x_{t})-f^{*}\geq 0.$
In summary, we have
$f^{*}\leq\liminf_{t\to\infty}f(x_{t})\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*},$
so strong duality holds. ∎
Unfortunately, 1 on its own is not enough to guarantee that Eq. 8 holds.
Example 2.
Define $X=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1\}\subset{\mathbb{R}}^{3}$, and $g(x)=x_{3}^{2}$. Then $g^{*}=0$ and $X_{g}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1,x_{3}=0\}$. Notice that for any $x\in X$, $\operatorname{Dist}(x,X_{g})=|x_{3}|=\sqrt{g(x)}$, therefore 1 is satisfied. Now define $f(x)=x_{2}^{2}/x_{1}$, which is convex. Then $f^{*}=0$ and $X^{*}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,x_{2}=0,x_{3}=0\}$. Now consider the sequence $\{x_{t}\}_{t\geq 1}$ defined by $x_{t}=(t,1,0)\in X_{g}$. Then $f(x_{t})=1/t\to f^{*}$, $g(x_{t})=0=g^{*}$, $\operatorname{Dist}(x_{t},X_{g})=0$ but $\operatorname{Dist}(x_{t},X^{*})=1$ for all $t$. Therefore Eq. 8 is not satisfied.
2 is more powerful than 1, and in fact has strong connections to the study of well-posedness, which examines properties of approximation schemes for solutions of optimization (and related) problems.
Definition 2 (Huang and Yang [15]).
We say that Eq. 2 satisfies:
  • Levitin-Polyak (LP) well-posedness if, for any sequence $\{x_{t}\}_{t\geq 1}\subset X$ which satisfies $\operatorname{Dist}(x,X_{g})\to 0$ and $f(x_{t})\to f^{*}$, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
  • generalized LP well-posedness if, for any sequence that satisfies Eq. 7, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
  • strong generalized LP well-posedness if, for any sequence that satisfies Eq. 5, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
It is immediately clear that strong generalized LP well-posedness implies generalized LP well-posedness. We will show that every sense of well-posedness in Definition 2 implies 2. It turns out that strong generalized LP well-posedness is equivalent to 2 in finite dimensions (which is our setting), which follows from a result of Huang and Yang [15]. We will additionally show that 2 ensures that strong duality holds, and thus Theorem 1 implies that generalized LP is equivalent to strong generalized LP well-posedness for Eq. 2.
Proposition 2.
The following are equivalent for Eq. 2:
  1. (a)
    2 holds.
  2. (b)
    Eq. 2 is generalized LP well-posed.
  3. (c)
    Eq. 2 is strongly generalized LP well-posed.
Furthermore, if Eq. 2 is LP well-posed, then 2 holds.
Proof of Proposition 2.
First, note that if $X^{*}$ is not compact, then it must be unbounded, since we assume that $X^{*}$ is closed. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X^{*}$ such that $\|x_{t}\|_{2}\to\infty$. Then this sequence satisfies all three premises of Definition 2, however there is no convergent subsequence. Therefore all three senses of well-posedness imply compactness of $X^{*}$, i.e., 2. Huang and Yang [15, Theorem 2.4] states that when $X$ is finite-dimensional, which we assume, then strong generalized LP well-posedness is equivalent to 2.
To show that strong generalized LP is equivalent to generalized LP, we will first show that under 2, strong duality holds. First, denote
$X_{1}:=\left\{x\in X:f(x)-f^{*}\leq 1,g(x)-g^{*}\leq 1\right\}.$
Note that $X_{1}$ is the $1$-sublevel set of the function $\bar{h}=\max\{f-f^{*},g-g^{*}\}$, and $X^{*}$ is the 0-sublevel set. Since $X^{*}$ is bounded, all sublevel sets of $\bar{h}$ are bounded, hence $X_{1}$ is also bounded. Let $\{x_{t}\}_{t\geq 1}\subset X$ be a sequence satisfying the premise of Lemma 1, i.e., $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. By weak duality $d^{*}\leq f^{*}$, therefore $f(x_{t})\leq f^{*}+1$ eventually, and also $g(x_{t})\leq g^{*}+1$ eventually as well. This means that only finitely many $x_{t}\not\in X_{1}$, and since $X_{1}$ is bounded, the whole sequence is bounded. Take a convergent subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$. Since $X$ is closed, $\bar{x}\in X$. Furthermore, since $g$ is convex, it is lower semicontinuous, so $g^{*}=\liminf_{k\to\infty}g(x_{t_{k}})\geq g(\bar{x})\geq g^{*}$, hence $\bar{x}\in X_{g}$. But this implies that $f(\bar{x})\geq f^{*}$ by definition of $f^{*}$. Therefore $f^{*}\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*}$, and thus strong duality holds.
Now, assume that Eq. 2 is generalized LP well-posed. Then $X^{*}$ must be compact. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfying Eq. 5. Since compactness of $X^{*}$ implies strong duality, Theorem 1, we know that $\{x_{t}\}_{t\geq 1}$ also satisfies Eq. 7. Therefore by generalized LP well-posedness, there exists $\bar{x}\in X^{*}$ and a subsequence for which $x_{t_{k}}\to\bar{x}$, hence Eq. 2 is also strongly generalized LP well-posed. Since a strongly generalized LP well-posed problem is also generalized LP well-posed, this shows the equivalence. ∎
We now use this equivalence to provide a quantitative version of Eq. 8.
Corollary 1.
Suppose that 2 holds for Eq. 2. Then there exists a function $c:{\mathbb{R}}_{+}^{2}\to{\mathbb{R}}_{+}$ such that $c(0,0)=0$ and for any $x\in X$,
$|f(x)-f^{*}|\geq c\left(\operatorname{Dist}(x,X^{*}),g(x)-g^{*}\right).$
Furthermore, the function $c$ satisfies that whenever $\beta_{t}\to 0$, $c(\alpha_{t},\beta_{t})\to 0$, then $\alpha_{t}\to 0$. Therefore, under 2 any sequence $\{x_{t}\}_{t\geq 1}$ satisfying Eq. 5 also satisfies Eq. 8.
Proof of Corollary 1.
Proposition 2 states that 2 is equivalent to generalized LP well-posedness. Huang and Yang [15, Theorem 2.2] then ensures that whenever Eq. 2 is generalized LP well-posed, the function $c$ that satisfies the required properties is
$c(\alpha,\beta):=\inf_{x}\left\{|f(x)-f^{*}|:x\in X,\operatorname{Dist}(x,X^{*})=\alpha,g(x)-g^{*}=\beta\right\}.$
4 Conclusion
We studied convergence guarantees for the convex bilevel problem (2). Through Example 1 we showed that typical bounds (5) are insufficient: they can hold while neither the objective values nor the iterates approach a bilevel solution in a meaningful sense. Our main result establishes an exact remedy: strong duality for the value-function formulation (3) is equivalent to the property that every sequence satisfying (5) necessarily enjoys objective convergence (7) (Theorem 1).
We then identified two simple conditions ensuring this regularity. First, the proximity assumption 1 guarantees strong duality and hence (7); however, as Example 2 demonstrates, it does not ensure convergence to the optimal solution set (8). Second, the compactness assumption 2 is stronger but widely utilized in previous work. It yields strong duality of Eq. 2 and solution-set convergence (Proposition 2) together with quantifiable bounds. These findings provide algorithm-agnostic criteria under which iterate sequences for Eq. 2 have meaningful limits.
Interestingly, Proposition 2 and Example 2 shows that Eq. 2 may still be ill-posed (in the sense of Definition 2) even when strong duality holds. Future work includes investigating the implications of the stronger notion of dual solvability on approximating sequences for Eq. 2. Furthermore, obtaining exact forms of the quantified bound in Corollary 1 for important bilevel problems is of interest.